# Video Compression Commander: Plug-and-Play Inference Acceleration for Video Large Language Models

### v1

@@ 1 Introduction
Recently, Video Large Language Models (VideoLLMs) have demonstrated remarkable performance in video understanding and reasoning tasks (Zhang et al., 2023; Wang et al., 2025). However, videos inherently contain multiple consecutive frames, resulting in a significantly higher number of visual tokens compared to images. For instance, LLaVA-OneVision (Li et al., 2024a) processes $32\times 196$ visual tokens per video, while LLaVA-Video (Zhang et al., 2024c) handles even more at $64\times 182$ visual tokens. This high token count inevitably leads to expensive computation during model inference, especially for long video understanding (Chen et al., 2024b).
To mitigate this computational burden, researchers have turned to token compression methods (Chen et al., 2024a; Yang et al., 2025), considering the inherent visual redundancy and aiming to minimize redundant visual information. These approaches can be categorized as pre-LLM (Zhang et al., 2024b) or intra-LLM (Chen et al., 2024a) methods, based on whether compression occurs before or within the LLM. Most of these methods are training-free, enabling plug-and-play inference acceleration for existing VideoLLMs. However, despite these efforts, existing token compression methods suffer from two critical issues:
(I) Design Myopia: In human video perception, we naturally focus on distinctive frames (e.g., those with significant spatio-temporal changes) while ignoring repetitive and redundant visual information. By contrast, most existing token compression methods apply a uniform compression strategy across all frames, treating each one as equally informative. Even recent VideoLLM-specific method DyCoke (Tao et al., 2025) exhibits this limitation by grouping every four consecutive frames into a fixed window and compressing them identically, without regard for the varying distinctiveness of individual frames. Figure 1 further illustrates the critical nature of this issue: removing 24 redundant frames does not affect the accurate response of the LLaVA-OneVision, whereas dropping just 8 unique frames causes it to fail, despite being only a third of the number. This contrast shows that uniform compression risks discarding critical information in unique frames that VideoLLMs may rely on, thereby significantly impacting overall performance. Notably, Table 2 indicates that some methods even underperform random token dropping, further indicating their sub-optimal performance.
@@
(II) Implementation Constraints: Beyond design limitations, existing methods face practical constraints. Some token compression works (Zhang et al., 2024b; Liu et al., 2024) rely on [CLS] attention weights in ViT for informative token preservation, yet modern VideoLLMs adopt SigLIP (Zhai et al., 2023) as visual encoder without [CLS] token. Meanwhile, certain methods (Zhang et al., 2025; Xing et al., 2025) aim to leverage textual information but require explicit attention weights in specific LLM layers, making them incompatible with efficient attention operators (Dao et al., 2022). This incompatibility leads to higher peak memory usage, even surpassing that of uncompressed processing (see Table 3), which is especially problematic for long video understanding (Wen et al., 2025).
We summarize existing works in Table 1 and identify three key principles for designing effective and efficient token compression methods for VideoLLM: (i) Model Adaptability: The method should be easily compatible with and adaptable to the majority of existing VideoLLMs (Zhang et al., 2024c; Wang et al., 2024); (ii) Frame Uniqueness: The method should consider varying distinctiveness across video frames; (iii) Operator Compatibility: The method should maintain compatibility with efficient operators (Dao, 2024).
@@ 2.1 Video Large Language Models
Large vision-language models (LVLMs) combine vision encoders with LLMs for exceptional visual understanding (Li et al., 2024a; Wang et al., 2024). While LVLMs can handle basic video tasks, the growing demand has led to specialized video large language models (VideoLLMs) (Zhang et al., 2024c, 2023). These VideoLLMs enhance video understanding through extensive datasets and targeted training strategies, as demonstrated by LLaVA-OneVision (Li et al., 2024a) for multi-modal tasks and LLaVA-Video (Zhang et al., 2024c) for video instruction-following. However, the long sequences of visual tokens from continuous video frames limit their practical applications.
@@ 2.2 Token Compression for LVLMs
Recently, with the increase in visual tokens in LVLMs, research has shifted from training-aware (Li et al., 2024c) to training-free token compression methods (Yang et al., 2025). Training-free approaches are generally categorized as: (a) Pre-LLM token compression at the ViT or projector level (Zhang et al., 2024b; Liu et al., 2025); (b) Intra-LLM token compression within the LLM decoder (Chen et al., 2024a; Zhang et al., 2025); and (c) Hybrid token compression that compresses tokens at both ViT and LLM (Liu et al., 2024; Han et al., 2024). However, these methods treat video frames as separate images, overlooking temporal relationships. While recent work DyCoke (Tao et al., 2025) introduces temporal token merging across consecutive frame windows, it cannot achieve retention ratios below 25%. More importantly, existing methods, including DyCoke, adopt uniform compression across frames without considering frame uniqueness, and many face compatibility issues with efficient operators (Dao et al., 2022).
@@ 3.1 Preliminary
Most current VideoLLMs follow the “ViT-MLP-LLM” paradigm (Li et al., 2024a; Zhang et al., 2024c). For example, in LLaVA-Video, a video sequence $\mathbf{V}=\left\{\mathbf{v}_{t}\right\}_{t=1}^{T}\in\mathbb{R}^{T\times H\times W\times 3}$ is first encoded by ViT into embeddings $\mathbf{Z}=\left\{\mathbf{z}_{t}\right\}_{t=1}^{T}\in\mathbb{R}^{T\times N\times D}$. These embeddings are projected by a 2-layer MLP and pooled to produce visual tokens $\mathbf{X}^{v}=\left\{\mathbf{x}_{t}^{v}\right\}_{t=1}^{T}\in\mathbb{R}^{T\times M\times D^{\prime}}$, with $M<N$, which are then fed into the LLM for autoregressive instruction-following:
@@
where $\bm{\Phi}$ represents the token compression operator and ||$\cdot$|| ⋅ | denotes the token length.
Token compression is particularly crucial for VideoLLMs due to their processing of substantially more visual tokens compared to standard LVLMs, a result of the multi-frame nature of videos. Consecutive frames often share high similarity, leading to significant visual redundancy. While recent method DyCoke (Tao et al., 2025) address some aspects of multi-frame redundancy, it struggles with uneven frame distinctiveness and achieving aggressive compression rates. Our work focuses on designing an effective token compression operator $\bm{\Phi}$ that adaptively handles frame-wise distinctiveness while enabling flexible compression rates, addressing these key challenges for VideoLLMs.
@@ 3.2 Video Compression Commander
To improve the computational efficiency of VideoLLMs, we propose “Video Compression Commander” (VidCom2), a novel token compression framework that adaptively minimizes visual redundancy within a predefined token budget while preserving distinctive visual information. VidCom2 maintains compatibility with efficient attention operators (Dao et al., 2022; Dao, 2024) and supports flexible compression rates, enabling plug-and-play inference acceleration.
@@ 3.3 Stage 1: Frame Compression Adjustment
where $\mathbf{g_{v}}$ serves as a coarse-grained summary of the entire video. Then, inspired by existing efforts (Sun et al., 2025), we compute the similarity between each token $\mathbf{x}_{t,m}^{v}$ and global video representation $\mathbf{g_{v}}$ in high-dimensional space:
@@ 4.1 Experimental Setting
Benchmark. We conduct comprehensive comparative experiments across multiple benchmarks, including: MVBench (Li et al., 2024b), LongVideoBench (Wu et al., 2024), MLVU (Zhou et al., 2024), and VideoMME (Fu et al., 2024), employing LMMs-Eval (Zhang et al., 2024a) evaluation framework. Details are in Appendix A.
Implementations. We evaluate our method on popular VideoLLMs: LLaVA-OneVision (LLaVA-OV) (Li et al., 2024a), LLaVA-Video (Zhang et al., 2024c), and Qwen2-VL (Wang et al., 2024). Detailed model information is in Appendix B. All experiments use NVIDIA A100-SXM4-80GB GPUs.
Baselines. We evaluate our method against various training-free token compression strategies, including: FastV (Chen et al., 2024a), PDrop (Xing et al., 2025), SparseVLM (Zhang et al., 2025), and DyCoke (Tao et al., 2025), more introduction can be seen in Appendix C. Following SparseVLM, we use the “equivalent retention ratio”111“Equivalent retention ratio” represents the average percentage of visual tokens retained across all LLM layers. for fair comparisons. Unlike others, DyCoke compresses both visual tokens and KV cache. For fair comparison, we evaluate only on its token compression strategy.
@@ 4.2 Main Comparisons
Figure 4: Performance on Qwen2-VL. At $R=25\%$, VidCom2 surpasses DyCoke and SparseVLM by 7.6% and 4.6% of original performance in long video tasks.
@@
Random 178.2 (↓71.2%) 566.0 (↓43.9%) 18:44 (↓28.1%) 16.0 (↓9.6%) 0.89 (1.39$\times$) 54.6 (↓2.3)
FastV [ECCV’24] 260.9 (↓57.8%) 648.6 (↓35.7%) 20:07 (↓22.8%) 24.7 (↑39.5%) 0.83 (1.30$\times$) 55.5 (↓1.4)
PDrop [CVPR’25] 205.6 (↓66.7%) 592.6 (↓41.2%) 18:50 (↓27.7%) 24.5 (↑38.4%) 0.88 (1.38$\times$) 55.3 (↓1.6)
SparseVLM [ICML’25] 410.6 (↓33.6%) 807.7 (↓19.9%) 25:03 (↓3.8%) 27.1 (↑53.1%) 0.67 (1.05$\times$) 56.4 (↓0.5)
DyCoke [CVPR’25] 205.2 (↓66.8%) 598.0 (↓40.7%) 18:56 (↓27.4%) 16.1 (↓9.0%) 0.88 (1.38$\times$) 49.5 (↓7.4)
VidCom2 180.7 (↓70.8%) 574.7 (↓43.0%) 18:46 (↓28.0%) 16.0 (↓9.6%) 0.88 (1.38$\times$) 57.2 (↑0.3)
@@
Random 130.3 (↓78.9%) 532.5 (↓47.2%) 18:02 (↓30.8%) 15.8 (↓10.7%) 0.92 (1.44$\times$) 53.1 (↓3.8)
FastV [ECCV’24] 172.4 (↓72.1%) 599.3 (↓40.6%) 18:19 (↓29.7%) 24.6 (↑39.0%) 0.91 (1.42$\times$) 51.6 (↓5.3)
PDrop [CVPR’25] 165.3 (↓73.3%) 552.6 (↓45.2%) 18:32 (↓28.9%) 24.5 (↑38.4%) 0.90 (1.41$\times$) 53.2 (↓3.7)
SparseVLM [ICML’25] 370.4 (↓40.1%) 764.8 (↓24.2%) 24:09 (↓7.3%) 27.1 (↑53.1%) 0.69 (1.08$\times$) 52.9 (↓4.0)
VidCom2 129.2 (↓79.1%) 533.0 (↓47.1%) 18:11 (↓30.2%) 15.8 (↓10.7%) 0.92 (1.44$\times$) 54.3 (↓2.6)
@@
(i) State-of-the-art Performance: VidCom2 demonstrates remarkable performance across various video understanding benchmarks. On LLaVA-OV and LLaVA-Video at $R=25\%$, VidCom2 surpasses DyCoke by significant margins of 12.6% and 7.3% respectively. Notably, VidCom2 with $R=25\%$ (99.6%) even outperforms DyCoke at higher compression rate $R=30\%$ (96.5%). This superiority extends to long video understanding with Qwen2-VL (Figure 4), where VidCom2 achieves 101.2% performance on VideoMME (Long), surpassing DyCoke (93.6%) and SparseVLM (96.6%) by 7.6% and 4.6%.
@@
Beyond performance, Table 3 presents comprehensive real-world inference efficiency comparisons among different token compression methods on MVBench, with all experiments conducted on four NVIDIA A100 GPUs. We follow the original implementation of each baseline method, and unless otherwise specified, Flash Attention 2 (Dao, 2024) is used as the efficient attention operator throughout comparisons. The comparison results in Table 3 reveal two key efficiency advantages of our VidCom2:
(i) State-of-the-art Efficiency: VidCom2 achieves remarkable inference efficiency, comparable to simple random token dropping. With 25% visual tokens retained, the additional computation of VidCom2 is negligible – only 2.5s extra (1.3% of LLM generation time) for the entire MVBench inference. Despite this minimal overhead, VidCom2 significantly reduces both the LLM generation latency and overall model latency (primarily from ViT and LLM) by 70.8% and 43.0% respectively, achieving 1.38$\times$ throughput while maintaining 99.6% average performance across benchmarks. These results highlight the efficiency of VidCom2 in accelerating inference for VideoLLMs.
@@ 4.3 Ablation Study and Analysis
For frame-level uniqueness, defining $u^{\mathrm{frame}}_{t,m}$ as the negative similarity to frame-level global representation ($-s^{\mathrm{frame}}_{t,m}$) outperforms positive similarity. Similarly, for video-level uniqueness, tokens less similar to the video-level global representation prove more informative. These results indicate that unique tokens, both within frames and across the video, should be prioritized during token compression to preserve richer visual information.
@@
Table 5: Effects of different compression adjustment. “Uniform”: fixed $R=25\%$. “$\max u^{\mathrm{video}}_{t,m}$” and “$\overline{u^{\mathrm{video}}_{t,m}}$” denote frame uniqueness score $u_{t}$ of frame $t$ computed by maximum and average operations of $u^{\mathrm{video}}_{t,m}$.
@@
Table 5 compares different compression adjustment strategies: (a) “Uniform” with fixed $R=25\%$ (no adjustment); (b)$\max\limits_{m}u^{\mathrm{video}}_{t,m}$” and (c)$\overline{u^{\mathrm{video}}_{t,m}}$”, which compute frame uniqueness score $u_{t}$ for token budget allocation using maximum and average operations of $u^{\mathrm{video}}_{t,m}$ in frame $t$, where larger $u_{t}$ leads to more tokens preserved in frame $t$.

### v2

@@ 1 Introduction
Recently, Video Large Language Models (VideoLLMs) have demonstrated remarkable performance in video understanding and reasoning tasks Zhang et al. (2023); Wang et al. (2025). However, videos inherently contain multiple consecutive frames, resulting in a significantly higher number of visual tokens compared to images. For instance, LLaVA-OneVision Li et al. (2024a) processes $32\times 196$ visual tokens per video, while LLaVA-Video Zhang et al. (2024c) handles even more at $64\times 182$ visual tokens. This high token count inevitably leads to expensive computation Liu et al. (2025b), especially for long video understanding Chen et al. (2024b).
To mitigate this computational burden, researchers have turned to token compression methods Chen et al. (2024a); Yang et al. (2025), considering the inherent visual redundancy and aiming to minimize redundant visual information. These approaches can be categorized as pre-LLM Zhang et al. (2024b) or intra-LLM Chen et al. (2024a) methods, based on whether compression occurs before or within the LLM. Most of these methods are training-free, enabling plug-and-play inference acceleration for existing VideoLLMs. However, despite these efforts, existing token compression methods suffer from two critical issues:
(I) Design Myopia: In human video perception, we naturally focus on distinctive frames (e.g., those with significant spatio-temporal changes) while ignoring repetitive and redundant visual information Ma et al. (2025). By contrast, most existing token compression methods apply a uniform compression strategy across all frames, treating each one as equally informative. Even recent VideoLLM-specific method DyCoke Tao et al. (2025) exhibits this limitation by grouping every four consecutive frames into a fixed window and compressing them identically, without regard for the varying distinctiveness of individual frames. Figure 1 further illustrates the critical nature of this issue: removing 24 redundant frames does not affect the accurate response of the LLaVA-OneVision, whereas dropping just 8 unique frames causes it to fail, despite being only a third of the number. This contrast shows that uniform compression risks discarding critical information in unique frames that VideoLLMs may rely on, thereby significantly impacting overall performance. Notably, Table 2 indicates that some methods even underperform random token dropping, further indicating their sub-optimal performance.
@@
(II) Implementation Constraints: Beyond design limitations, existing methods face practical constraints. Some token compression works Zhang et al. (2024b); Liu et al. (2024) rely on [CLS] attention weights in ViT for informative token preservation, yet modern VideoLLMs adopt SigLIP Zhai et al. (2023) as visual encoder without [CLS] token. Meanwhile, certain methods Zhang et al. (2025); Xing et al. (2025) aim to leverage textual information but require explicit attention weights in specific LLM layers, making them incompatible with efficient attention operators Dao et al. (2022). This incompatibility leads to higher peak memory usage, even surpassing that of uncompressed processing (see Table 4), which is especially problematic for long video understanding Wen et al. (2025a, b).
We summarize existing works in Table 1 and identify three key principles for designing effective and efficient token compression methods for VideoLLM: (i) Model Adaptability: The method should be easily compatible with and adaptable to the majority of existing VideoLLMs Zhang et al. (2024c); Wang et al. (2024); (ii) Frame Uniqueness: The method should consider varying distinctiveness across video frames; (iii) Operator Compatibility: The method should maintain compatibility with efficient operators Dao (2024).
@@ 2.1 Video Large Language Models
Large vision-language models (LVLMs) combine vision encoders with LLMs for exceptional visual understanding Li et al. (2024a); Wang et al. (2024). While LVLMs can handle basic video tasks, the growing demand has led to specialized video large language models (VideoLLMs) Zhang et al. (2024c, 2023). These VideoLLMs enhance video understanding through extensive datasets and targeted training strategies, as demonstrated by LLaVA-OneVision Li et al. (2024a) for multi-modal tasks and LLaVA-Video Zhang et al. (2024c) for video instruction-following. However, the long sequences of visual tokens from continuous video frames limit their practical applications.
@@ 2.2 Token Compression for LVLMs
Recently, with the increase in visual tokens in LVLMs, research has shifted from training-aware Li et al. (2024c) to training-free token compression methods Yang et al. (2025). Training-free approaches are generally categorized as: (a) Pre-LLM token compression at the ViT or projector level Zhang et al. (2024b); Liu et al. (2025a); (b) Intra-LLM token compression within the LLM decoder Chen et al. (2024a); Zhang et al. (2025); Chen et al. (2025); and (c) Hybrid token compression that compresses tokens at both ViT and LLM Han et al. (2024). However, these methods treat video frames as separate images, overlooking temporal relationships. While recent work DyCoke Tao et al. (2025) introduces temporal token merging across consecutive frame windows, it cannot achieve retention ratios below 25%. More importantly, existing methods, including DyCoke, adopt uniform compression across frames without considering frame uniqueness, and many face compatibility issues with efficient operators Dao et al. (2022).
@@ 3.1 Preliminary
Most current VideoLLMs follow the “ViT-MLP-LLM” paradigm Li et al. (2024a); Zhang et al. (2024c). For example, in LLaVA-Video, a video sequence $\mathbf{V}=\left\{\mathbf{v}_{t}\right\}_{t=1}^{T}\in\mathbb{R}^{T\times H\times W\times 3}$ is first encoded by ViT into embeddings $\mathbf{Z}=\left\{\mathbf{z}_{t}\right\}_{t=1}^{T}\in\mathbb{R}^{T\times N\times D}$. These embeddings are projected by a 2-layer MLP and pooled to produce visual tokens $\mathbf{X}^{v}=\left\{\mathbf{x}_{t}^{v}\right\}_{t=1}^{T}\in\mathbb{R}^{T\times M\times D^{\prime}}$, with $M<N$, which are then fed into the LLM for autoregressive instruction-following:
@@
where $\bm{\Phi}$ represents the token compression operator and $|\cdot|$ denotes the token length.
Token compression is particularly crucial for VideoLLMs due to their processing of substantially more visual tokens compared to standard LVLMs, a result of the multi-frame nature of videos. Consecutive frames often share high similarity, leading to significant visual redundancy. While recent method DyCoke Tao et al. (2025) address some aspects of multi-frame redundancy, it struggles with uneven frame distinctiveness and achieving aggressive compression rates. Our work focuses on designing an effective token compression operator $\bm{\Phi}$ that adaptively handles frame-wise distinctiveness while enabling flexible compression rates, addressing these key challenges for VideoLLMs.
@@ 3.2 Video Compression Commander
To improve the computational efficiency of VideoLLMs, we propose “Video Compression Commander” (VidCom2), a novel token compression framework that adaptively minimizes visual redundancy within a predefined token budget while preserving distinctive visual information. VidCom2 maintains compatibility with efficient attention operators Dao et al. (2022); Dao (2024) and supports flexible compression rates, enabling plug-and-play inference acceleration.
@@ 3.3 Stage 1: Frame Compression Adjustment
where $\mathbf{g_{v}}$ serves as a coarse-grained summary of the entire video. Then, inspired by existing efforts Sun et al. (2025), we compute the similarity between each token $\mathbf{x}_{t,m}^{v}$ and global video representation $\mathbf{g_{v}}$ in high-dimensional space:
@@ 4.1 Experimental Setting
Benchmark. We conduct comprehensive comparative experiments across multiple benchmarks, including: MVBench Li et al. (2024b), LongVideoBench Wu et al. (2024), MLVU Zhou et al. (2024), VideoMME Fu et al. (2024), EgoSchema Mangalam et al. (2023), and PerceptionTest Patraucean et al. (2023), employing LMMs-Eval Zhang et al. (2024a) evaluation framework. More details are in Appendix A.
Implementations. We evaluate our method on popular VideoLLMs: LLaVA-OneVision (LLaVA-OV) Li et al. (2024a), LLaVA-Video Zhang et al. (2024c), and Qwen2-VL Wang et al. (2024). Detailed model information is in Appendix B. All experiments use NVIDIA A100-SXM4-80GB GPUs.
Baselines. We evaluate our method against various training-free token compression strategies, including: FastV Chen et al. (2024a), PDrop Xing et al. (2025), SparseVLM Zhang et al. (2025), and DyCoke Tao et al. (2025), more introduction can be seen in Appendix C. Following SparseVLM, we use the “equivalent retention ratio”111“Equivalent retention ratio” represents the average percentage of visual tokens retained across all LLM layers. for fair comparisons. Unlike others, DyCoke compresses both visual tokens and KV cache. For fair comparison, we evaluate only on its token compression strategy.
Figure 4: Performance with Qwen2-VL. At $R=25\%$, VidCom2 surpasses DyCoke and SparseVLM by 7.6% and 4.6% of original performance in long video tasks.
@@ 4.2 Main Comparisons
Methods EgoSchema PerceptionTest
Upper Bound
LLaVA-OV-7B 60.4 (100%) 57.1 (100%)
Retention Ratio=25%
FastV [ECCV’24] 57.5 (95.2%) 55.4 (97.0%)
PDrop [CVPR’25] 58.0 (96.0%) 55.6 (97.4%)
DyCoke [CVPR’25] 59.5 (98.5%) 56.4 (98.8%)
VidCom2 59.7 (98.8%) 56.7 (99.3%)
Table 3: Performance comparison on EgoSchema and PerceptionTest. Percentages represent ratios to the original performance of LLaVA-OV-7B.
@@
[…]
@@
[…]
@@
(i) State-of-the-art Performance: VidCom2 demonstrates exceptional performance across diverse video understanding benchmarks. On LLaVA-OV and LLaVA-Video with compression ratio $R=25\%$, VidCom2 substantially outperforms DyCoke by margins of 12.6% and 7.3%, respectively. Remarkably, VidCom2 at $R=25\%$ (achieving 99.6% performance retention) even surpasses DyCoke operating at a higher compression ratio of $R=30\%$ (96.5% performance retention). This superiority extends to long-form video understanding tasks with Qwen2-VL (Figure 4), where VidCom2 achieves 101.2% performance on VideoMME (Long), surpassing both DyCoke (93.6%) and SparseVLM (96.6%) by substantial margins of 7.6% and 4.6%, respectively. Additional comparisons in Table 3 further validate the superior performance advantages of VidCom2 across various video understanding scenarios.
@@
Beyond performance, Table 4 presents comprehensive real-world inference efficiency comparisons among different token compression methods on MVBench, with all experiments conducted on four NVIDIA A100 GPUs. We follow the original implementation of each baseline method, and unless otherwise specified, Flash Attention 2 Dao (2024) is used as the efficient attention operator throughout comparisons. The comparison results in Table 4 reveal two key efficiency advantages of our VidCom2:
[…]
@@ 4.3 Ablation Study and Analysis
For frame-level uniqueness, defining $u^{\mathrm{frame}}_{t,m}$ as the negative similarity to frame-level global representation ($s^{\mathrm{frame}}_{t,m}$) outperforms positive similarity. Similarly, for video-level uniqueness, tokens less similar to the video-level global representation prove more informative. These results indicate that unique tokens, both within frames and across the video, should be prioritized during token compression to preserve richer visual information.
@@
Table 6: Effects of different compression adjustment. “Uniform”: fixed $R=25\%$. $\max u^{\mathrm{video}}_{t,m}$” and $\overline{u^{\mathrm{video}}_{t,m}}$” denote frame uniqueness score $u_{t}$ of frame $t$ computed by maximum and average operations of $u^{\mathrm{video}}_{t,m}$.
@@
[…]
@@
Size MVBench VideoMME Avg.
Overall Short Medium Long
Vanilla 56.9 58.6 70.3 56.6 48.8 100.0
4 56.8 57.9 69.6 55.6 48.7 99.1
8 56.8 58.3 69.8 56.4 48.6 99.6
16 57.2 58.5 70.0 56.7 48.9 100.1
32 57.2 58.6 69.8 56.4 49.4 100.1
Table 7: Effects of different window sizes for local $g_{v}$ computation. Window sizes up to 32 (global perspective) are evaluated on LLaVA-OV-7B.
Effects of Different Window Sizes for Local $g_{v}$ Computation
We explore sliding window strategies for computing local $g_{v}$ representations to investigate the effectiveness of adjusting frame compression intensity from local perspectives. We evaluate different window sizes (4, 8, 16, 32) on LLaVA-OV-7B with fixed 32 frames.
As shown in Table 7, performance consistently improves as window size increases across both MVBench and VideoMME. Notably, when window sizes reach 16 and 32, the performance gap becomes marginal. Window size 16 achieves better results on VideoMME short and medium videos, while window size 32 (global perspective) demonstrates superior performance on VideoMME long videos. Therefore, we adopt the global perspective for adjusting compression intensity by default to achieve better long video understanding.